OpenStudy (yellowlegoguy99):

Simplify 3^-9 / 3^-12 1 / 3^3 1 / 3^21 3^3 3^21

OpenStudy (jack1):

go Happy Feet!

OpenStudy (jack1):

like... u go girl...?

OpenStudy (jack1):

cheering...?

OpenStudy (loser66):

hahahaha... you mean go to see girl?

OpenStudy (jack1):

no, i know what i meant... like, answer the question, and I'll cheer u on from the sidelines

OpenStudy (yellowlegoguy99):

......

OpenStudy (jack1):

coz ur a caring person... it's in ur nature to care ;D

OpenStudy (yellowlegoguy99):

Can someone answer the question?

OpenStudy (loser66):

hi, yellow guy, let this blue tell you something, when making question, should stay to get help. Jack1 will help you, right, Jack?

OpenStudy (jack1):

aye lassie, np

OpenStudy (jack1):

so same as last time man 3^-9 / 3^-12 = 3^-9 x 3^12 = 3^(-9 + 12) = 3 ^(3)

OpenStudy (jack1):

= 27

OpenStudy (loser66):

good job, Jack hehehehe...

OpenStudy (jack1):

see, u do care @ penguin

OpenStudy (loser66):

I watch you out in case you give out the wrong answer hehahahaha... joking Jack!!

OpenStudy (yellowlegoguy99):

So the answer is C, right?

OpenStudy (jack1):

u should, im new here

OpenStudy (jack1):

yep, the divide by turns the -12 into +12 power add the powers together (same base) final answer = 3^3